which fraction is larger 1/3, 2/5, 8/15, 13/30. and how do i solve this
Found 2 solutions by longjonsilver, sdmmadam@yahoo.com:
Answer by longjonsilver(2297)

1/3, 2/5, 8/15, 13/30

we need to convert them (scale them) so that they have the same denominators.

10/30, 2/5, 8/15, 13/30
10/30, 12/30, 8/15, 13/30
10/30, 12/30, 16/30, 13/30

Now, they are all 30ths.... which is largest? Answer is 16/30 --> 8/15.

jon.

Answer by sdmmadam@yahoo.com(530)

which fraction is larger 1/3, 2/5, 8/15, 13/30. and how do i solve this
Whenever you are given numerical fractions and asked to find the largest,smallest etc. what is to be done is without changing the meaning bring them to a common denominator and then denominator being the same looking at the value of the nrs we can decide which is bigger than what fraction.
How to bring the drs to a common denominator. What we mean by a common denominator is that "same denominator"
Find the lcm of all the denominators of the given fractions convert every fraction to a fraction whose dr is this lcm value.
Now given 1/3, 2/5, 8/15, 13/30
the lcm of 3,5,15 and 30 is 30
and 3X(10)= 30, 5X(6) = 30, 15X(2) = 30
Therefore (1/3) = (1/3)X(10/10)=(10/30),
(2/5)= (2/5)X(6/6)=(12/30),
(8/15) = (8/15)X(2/2)=(16/30)
and therefore 1/3,2/5, 8/15, 13/30 become respectively
10/30,12/30,16/30 and 13/30
Now studying the nrs, 10,12,16,and 13, arranging them in ascending order (that is from smallest to the next bigger to the next bigger ... to the biggest)
We have 10,12,13 and 16
Therefore the fractions in the ascending order are 10/30,12/30,13/30 and 16/30
That is 1/3,2/5, 13/30 ,8/15
Therefore 1/3 < 2/5 < 13/30 < 8/15
(1/3) is the smallest
larger than (1/3) is (2/5)
larger than (2/5) is (13/30)
The largest is (8/15)
Note: make a neat copy of the above notes removing all explanatory text.That will be your answer for the test.